OpenStudy (anonymous):

find a linear function h such that h(3)=7 and h(-1)=14. What is h(1/2)?

OpenStudy (anonymous):

k so u must know the formula slope = y2-y1/x2-x1

OpenStudy (anonymous):

now plug in your coordinates in order 14-7/-1-3 = 7/-4

OpenStudy (anonymous):

k dunno wut happened there .... anyway then use one of the points for example (3,7) in the y = mx + b equation

OpenStudy (anonymous):

m = -7/4 , x= 3 and y = 7 .... 7 = (-7/4)3+b .... b = 59/4

OpenStudy (anonymous):

plug in x = 1/2 .... you get 111/8 ..... very ugly numbers i must say